OpenStudy (anonymous):

how to find the derivative of f(x) = 6x + 2 at x = 1.

OpenStudy (irishboy123):

if you are looking at how things change relatively - which is why derivatives are important - then: \(\dfrac{d}{dx} x^n = n \; x^{n-1}\) and constants don't change. so \(\dfrac{d}{dx} C= \; 0\) that could be pretty cryptic, or very obvious, so tag me if you need more input.